Chris says, I am thinking of two positive integers that have a difference of 8 and the sum of their squares is 610." Algebraically determine the integers that Chris is thinking of.
Answer by mananth(12270)

if one integer is x
the other will be x-8
x^2+(x-8)^2=610
x^2+x^2-16x+64=610
2x^2-16x=546
/2
x^2-8x=273
x^2-8x-273=0
x^2-21x+13x-273=0
x(x-21)+13(x-21)=0
(x-21)(x+13)=0
x=21 OR -13
The integers are positive
so the numbers are 21 & 13